OpenStudy (anonymous):

Identify the 16th term of a geometric sequence where a1 = 4 and a8 = −8,748. −172,186,884 −57,395,628 57,395,628 172,186,884

OpenStudy (cwrw238):

first find the common ratio r a8 / a1 = r^7 = -8748 / 4

OpenStudy (cwrw238):

16th term = a1 * r^15

OpenStudy (anonymous):

\[a_n=a_1*r^{n-1}\]\[a_8=a_1*r^{8-1}\]\[r^7=\frac{a_8}{a^1}\]\[r=\sqrt[7]{\frac{a_8}{a_1}}\]\[r=-3\]\[a_n=a_1*(-3)^{n-1}\]\[a_{16}=a_1*(-3)^{16-1}\]\[a_{16}=a_1*(-3)^{15}\]\[a_{16}=4*(-3)^{15}\]\[\boxed{a_{16}=-57,395,628}\]

OpenStudy (anonymous):

thank you both